OpenStudy (crisulcampo):

Good afternoon buddies, could you express this into ctg(alpha) terms? tg^2 (alpha) + tg(alpha) -tg(alpha)*sec^2 (alpha) thanks

OpenStudy (mathstudent55):

Use this identity for \(\tan \theta\) \(\tan \alpha = \dfrac{1}{\cot \alpha} \) We need to take care of \(\sec^2\theta\) \(1 + \tan^2 \theta = \sec^2 \theta\)

OpenStudy (crisulcampo):

mmmm perfect my friend, I already know that. However my math book answer says that the result is this : \[\frac{ 4 + 3ctan^2(\alpha) + 2ctan^4(\alpha) }{ctan^2(\alpha)[1 +ctan^2(\alpha)] }\] I got another different... So, could you get this expression in ctan(alpha) function factoring ?

OpenStudy (agent0smith):

\[\large \tan^2 \alpha + \tan \alpha -\tan \alpha \sec^2 \alpha\] \[\large \frac{ 1 }{ \cot^2 \alpha } + \frac{ 1 }{ \cot \alpha } - \frac{ 1 }{ \cot \alpha } (1 + \frac{ 1 }{ \cot^2 \alpha })\] \[\large \frac{ 1 }{ \cot^2 \alpha } + \frac{ 1 }{ \cot \alpha } - \frac{ 1 }{ \cot \alpha } + \frac{ 1 }{ \cot^3 \alpha }\] \[\large \frac{ 1 }{ \cot^2 \alpha } + \frac{ 1 }{ \cot^3 \alpha }= \frac{ \cot \alpha + 1 }{ \cot^3 \alpha}\]

OpenStudy (crisulcampo):

@agent0smith your procedure is wrong

OpenStudy (crisulcampo):

there's a little mistake

OpenStudy (agent0smith):

Makes very litle diff. though\[\large \frac{ 1 }{ \cot^2 \alpha } + \frac{ 1 }{ \cot \alpha } - \frac{ 1 }{ \cot \alpha } - \frac{ 1 }{ \cot^3 \alpha }\] \[\large \frac{ 1 }{ \cot^2 \alpha } - \frac{ 1 }{ \cot^3 \alpha }= \frac{ \cot \alpha - 1 }{ \cot^3 \alpha}\]

OpenStudy (crisulcampo):

that's the answer I got but, math book answer says it's the one I wrote before above

OpenStudy (crisulcampo):

I'm solving it... each one of the trygonometry exercises. I have discovered there are some mistakes

OpenStudy (agent0smith):

I was gonna try simplifying the book's expression, till i realized the numerator of theirs is prime, so I didn't even bother.